What is the relationship between standard deviation and accuracy?
It depends what you're asking. The question is extremely unclear. Accuracy of what exactly? Even in the realm of statistics an entire book could be written to address such an ambiguous question (to answer a myriad of possible questions).
If you simply are asking what the relationship between the probability that something will occur given the know distribution of outcomes (such as a normal distribution), the mean of that that distribution, and the the standard deviation, then the standard deviation as a represents the spread of the curve of probability. This means that if you had a cure where 0 was the mean, and 3 was the standard deviation, the likelihood of observing a value of 12 (or -12) would be likely inaccurate if that was your prediction. However, if you had a mean of 0 and a standard deviation of 100, the likelihood of observing of a 12 (or -12) would be quite likely. This is simply because the standard deviation provides a simple representation of the horizontal spread of probability on the x-axis.
What is sampling with replacement and why is it not used?
It is used. However, it can result in the same unit being observed again. That reduced the range of possible outcomes.
It is used. However, it can result in the same unit being observed again. That reduced the range of possible outcomes.
It is used. However, it can result in the same unit being observed again. That reduced the range of possible outcomes.
It is used. However, it can result in the same unit being observed again. That reduced the range of possible outcomes.
Is binomial probability distribution always negatively skewed?
No. It depends on the probability of success, p. If p < 0.5 the distribution is positively skewed.
No. It depends on the probability of success, p. If p < 0.5 the distribution is positively skewed.
No. It depends on the probability of success, p. If p < 0.5 the distribution is positively skewed.
No. It depends on the probability of success, p. If p < 0.5 the distribution is positively skewed.
What is the probability of a spinner stopping on a composite number when number 1-8 are shown?
It is 3/8
It is 3/8
It is 3/8
It is 3/8
What is a continuous random variable?
In the simplest setting, a continuous random variable is one that can assume any value on some interval of the real numbers.
For example, a uniform random variable is often defined on the unit interval [0,1], which means that this random variable could assume any value between 0 and 1, including 0 and 1. Some possibilities would be 1/3, 0.3214, pi/4, e/5, and so on ... in other words, any of the numbers in that interval.
As another example, a normal random variable can assume any value between -infinity and +infinity (another interval). Most of these values would be extremely unlikely to occur but they would be possible. The random variable could assume values of 3, -10000, pi, 1000*pi, e*e, ... any possible value in the real numbers.
It is also possible to define continue random variables that assume values on the entire (x,y) plane, or just on the circumference of a circle, or anywhere that you can imagine that is essentially equivalent (in some sense) to pieces of a real line.
The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.
The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.
The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.
The answer depends on what the trials are, what constitutes a favourable outcome, and how many possible outcomes there were previously. And since you have not bothered to provide these crucial bits of information, I cannot provide a more useful answer.
The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.
The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.
The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.
The answer depends on how many cards are drawn and whether or not they are replaced before the next card is drawn.
How many different ways can you arrange the letters of the word Malayalam?
There are 9!/[2!*4!*2!] = 3780 if you treat M and m as equivalent. Otherwise, there are twice as many.
What is probability compliment?
"That was a good bet!", perhaps.
"That was a good bet!", perhaps.
"That was a good bet!", perhaps.
"That was a good bet!", perhaps.
Skewness is 0, but kurtosis is -3, not 3.
Skewness is 0, but kurtosis is -3, not 3.
Skewness is 0, but kurtosis is -3, not 3.
Skewness is 0, but kurtosis is -3, not 3.
How many possible outcomes for normal distribution?
What is the event space for rolling a number greater than 2 with one die?
Why is calculating the standard deviation an important part of data analysis?
It is a measure of the spread of the results around their expected value.
It is a measure of the spread of the results around their expected value.
It is a measure of the spread of the results around their expected value.
It is a measure of the spread of the results around their expected value.
What proportion of a normal distribution is located between z-0.90 and z 0.90?
0.368 or 36.8%.
And you should specify that it is a standard normal distribution.
0.368 or 36.8%.
And you should specify that it is a standard normal distribution.
0.368 or 36.8%.
And you should specify that it is a standard normal distribution.
0.368 or 36.8%.
And you should specify that it is a standard normal distribution.
What is the probability of not geeting a queen in a well shuffled deck of 52 playing cards?
If only one card is selected the probability is 12/13.
If only one card is selected the probability is 12/13.
If only one card is selected the probability is 12/13.
If only one card is selected the probability is 12/13.
1316
What is the probability of drawing red cards 2 through 3?
The question does not say how many cards are drawn and so it is impossible to answer.
The question does not say how many cards are drawn and so it is impossible to answer.
The question does not say how many cards are drawn and so it is impossible to answer.
The question does not say how many cards are drawn and so it is impossible to answer.
What important question must you answer before computing an or probability?
You need to know whether or not the events are mutually exclusive.
What is the probability of flipping two coins and one landing on heads and one landing on tails?
It is 1/2.
What is the probability that Jen will roll a 3 on a number cube and flip tails on a coin?
There is a 1/6 chance for the 3 and a 1/2 chance for the coin. Multiply them to get a 1/12 chance for both. ~.0833%
What is the probability of obtaining at least one head when a coin is flipped four times?
The probability of obtaining 4 tails when a coin is flipped 4 times is:
P(4T) = (1/2)4 = 1/16 = 0.0625
Then, the probability of obtaining at least 1 head when a coin is flipped 4 times is:
P(at least 1 head) = 1 - 1/16 = 15/16 = 0.9375
When you role a die there 36 possible combinations What is the probability of each possible outcome?
Each outcome is equally likely and so the probability of each outcome is 1/36.